Inferno sulphuric acid H2 s o4 the oxidation number of sulphur is

Grade:12

1 Answers

Esha Mukherjee
5 years ago
Let's take oxidation state of sulphur as x
For oxygen it's -2 on each atom as there is no peroxide linkage
For hydrogen it's +1 for each atom
So +1(2)+x-2(4)=0
2+x-8=0
x-6=0
So x=6
Oxidation state of sulphur is +6